OpenStudy (anonymous):

Given f (x) = cos^2 (x) what is the correct expansion for the first derivative of f (x) from first principals:

OpenStudy (anonymous):

i think it is -sin2x

OpenStudy (anonymous):

perfect condition for me now, hihi..

OpenStudy (anonymous):

its wrong

OpenStudy (anonymous):

it is -2cosxsinx

OpenStudy (anonymous):

use the chain rule

OpenStudy (anonymous):

\[\cos ^{2}x = (\cos x)^{2}\] \[f \prime (x) = 2 \cos x \times -\sin x\]

OpenStudy (anonymous):

u need 2 use the first principal